how do i solve a
log with a base of 3 and exponent of 81

Answer by stanbon(48569)

log with a base of 3 and exponent of 81
---------------
log(base 3) 81 = x
3^x = 81
3^x = 3^4
x = 4
---------------
Cheers,
Stan H.